# A multiverse model in dS wedge holography

### v2

@@ 2.2 Multiverse models from the (JT) dS wedge holography
We now describe a multiverse model by gluing n-copies of dS wedge universes. The different UV and IR branes are pairwise glued together in the configuration illustrated in Fig. 2.
Figure 2: Penrose diagram of the AdS3 space (in gray) cutoff by UV and IR dS branes (in red). We prepare the state with the Hartle-Hawking construction, considering a thermal density matrix for each brane $\rho_{\rm TFD}$ purified by a maximally entangled partner (UV or IR brane respectively) with an empty bulk in between (white) and glue them together (red arrows), as shown in the figure. Although the UV and IR branes within a single universe are initially unentangled, as the branes accelerate and exchange Hawking modes.
@@
    $n$ double-sided IR and $n$ UV dSd branes with a matter theory entangled with each other.
@@
    $n$ double-sided AdSd+1 bulk space cutoff by ETW branes glued along pairs of UV/UV or IR/IR branes periodically.
In the simplest case, $d=2$, gravity would be topological. To generate a dynamical evolution with respect to a braneworld observer, we keep only the tension on the UV branes, and we add an intrinsic gravitational theory, namely dS JT gravity, on the IR brane. [^10] Alternatively, we could have added dS JT gravity on both the UV and IR branes. One can immediately see that adding JT on the UV brane would modify the phases for the entanglement in (34, 35) by the same additive (time-dependent) constant. Since only the relative entropy difference enters in the physical analysis, we do not consider this possibility, although it can be straightforwardly included. We now analyze the conditions for this configuration to be realized.
UV branes
The theory on each of the UV braneworlds is given by
$\int_{Q_{c}}\mathrm{d}^{2}x\sqrt{-h}\mathcal{L}_{\rm intrinsic}=\frac{1}{8\pi G_{3}}\int_{Q_{c}}\mathrm{d}^{2}x\sqrt{-h}K-\mathcal{T}\int_{Q_{c}}\mathrm{d}^{2}x\sqrt{-h}$ (8)
where $\mathcal{T}$ is the tension of the branes. The brane equation of motion translates to the Israel junction conditions Israel:1966rt
$\Delta K_{ij}-h_{ij}\Delta K=8\pi G_{3}T_{ij}$ (9)
where $T_{ij}$ is the effective stress tensor on the brane; while $\Delta K_{ij}=K_{ij}^{(+)}-K_{ij}^{(-)}=2K_{ij}^{(+)}$ is the extrinsic curvature difference between opposite sides, denoted by $\pm$, in the double-sided (i.e. the UV and UV, or IR and IR) $Q_{c}$ brane setting.
We consider an outward-directed normal vector $n^{\mu}$, at the location of the (+)(+)( + ) side of the UV brane $Q_{c}$. One finds
$K^{(+)}_{ij}=\frac{1}{2}[\partial_{n}h_{ij}]_{\sigma_{c}}=\coth\sigma_{c}\,h_{ij}$ (10)
where $\partial_{n}=\partial_{\sigma}$. The stress tensor from the brane contribution in (8) then determines the location in terms of the tension of the brane through the relation:
$\mathcal{T}=\frac{1}{4\pi G_{3}}\coth\sigma_{c}\,.$ (11)
@@
On the other hand, consider dS JT gravity as intrinsic theory on the IR branes, so that the system is described by
@@
The dilaton equation of motion fixes the background geometry on the ETW brane to be
@@
Remark on perturbative stability
Although the intrinsic gravity on the branes is decoupled from the ambient AdS3 gravity as the UV brane is placed at the conformal boundary, small fluctuations of the UV and IR branes could in principle be unstable in this system. To see this, notice that since the IR branes have an opposite orientation with respect to the UV branes, then the extrinsic curvature should be expressed as $K_{ij}^{(+)}=-\coth\sigma_{b}\leavevmode\nobreak\ h_{ij}$, which then implies in (9) that
$T_{ij}=-\frac{1}{4\pi G_{3}}\coth\sigma_{b}h_{ij}.$ (23)
Since $\sigma_{b}>0$ there would thus be negative components in the stress tensor on the brane, which can then lead to instabilities. In Aguilar-Gutierrez:2023tic a similar type of issue, when the IR brane acquires negative tension, was repaired by making an orbifold projection, such that the unstable mode is eliminated (see also Geng:2022slq ; Geng:2022tfc ). We repeat the same type of argument in our case. Consider the case of a multiverse with a single pair of double sided UV and IR branes as in Fig. 1 (the generalization with an arbitrary number of double sided branes follows the same consideration); the total action, including the dS JT gravity theory on the IR brane, is given by:
$I=$ $\frac{1}{8\pi G_{3}}\int_{\hat{\mathcal{M}}}\mathrm{d}^{3}x(\hat{R}+2)+\int_{Q_{c}}\mathrm{d}^{2}x\sqrt{-h}\quantity(\frac{1}{8\pi G_{3}}\Delta K-\mathcal{T})$ (24)
$+\frac{1}{8\pi G_{3}}\int_{Q_{b}}\mathrm{d}^{2}x\sqrt{-h}\Delta K+\frac{1}{8\pi G_{\rm b}}\int_{Q_{b}}\mathrm{d}^{2}x\sqrt{-h}\quantity[\Phi_{0}\tilde{R}+\Phi\quantity(\tilde{R}-2\Lambda_{\rm JT})]\leavevmode\nobreak\ .$
Notice the additional factors of 2 in the total action given that we are considering two-sided branes $Q_{c}$ and $Q_{b}$ as in Fig. 1. Now, we let the location of the branes fluctuate by allowing:
$\sigma_{b}\rightarrow\sigma_{b}+\delta\varphi_{b}\leavevmode\nobreak\ ,\quad\sigma_{c}\rightarrow\sigma_{c}+\delta\varphi_{c}\leavevmode\nobreak\ ,$ (25)
where $\delta\varphi_{c,\leavevmode\nobreak\ b}$ represent scalar mode fluctuations. Using the dS2 foliation of AdS3 space:
$\mathrm{d}s_{3}^{2}=\mathrm{d}\sigma^{2}+\sinh^{2}\sigma\quantity(g_{\mu\nu}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu})\leavevmode\nobreak\ ,$ (26)
we can express (24) with (25) as
$I_{\rm tot}=$ $\frac{1}{8\pi G_{3}}\int\mathrm{d}^{2}x\sqrt{-g}\int_{\sigma_{b}+\delta\varphi_{b}}^{\sigma_{c}+\delta\varphi_{c}}\mathrm{d}\sigma\quantity(R[g]-2\cosh 2\sigma)$ (27)
$+\frac{1}{4\pi G_{3}}\int\mathrm{d}^{2}x\sqrt{-g}\sinh^{2}(\sigma_{c}+\delta\varphi_{c})[2\coth(\sigma_{c}+\delta\varphi_{c})-\coth(\sigma_{c})]$
$-\frac{1}{4\pi G_{3}}\int\mathrm{d}^{2}x\sqrt{-g}\sinh^{2}(\sigma_{b}+\delta\varphi_{b})2\coth(\sigma_{b}+\delta\varphi_{b})$
[…]
where we have used ${h}_{ij}=\sinh^{2}\sigma\leavevmode\nobreak\ g_{ij}$, the dilaton equation of motion (14), and (15). Notice that $\Phi_{0}$ is a free parameter, and by fixing it to
$\Phi_{0}=\frac{G_{\rm b}}{G_{3}}\sinh\sigma_{b}\cosh\sigma_{b}\leavevmode\nobreak\ ,$ (28)
we can then express (27) as the same effective action describing scalar fluctuations as in Aguilar-Gutierrez:2023tic ; namely
$I_{\rm tot}=\frac{1}{8\pi G_{3}}\int\mathrm{d}^{2}x\sqrt{-g}\evaluated{\quantity[(\sigma_{i}+\delta\varphi_{i})R[g]-2\Lambda_{i}]}_{i=b}^{i=c}\leavevmode\nobreak\ ,$ (29)
where $\Lambda_{i}=\frac{1}{2}\sinh(2(\sigma_{i}+\delta\varphi_{i}))\quantity(\tfrac{\coth(\sigma_{i})}{\coth(\sigma_{i}+\delta\varphi_{i})}-1)$. The effective action to second order in perturbation theory is then
$\frac{1}{8\pi G_{3}}\int\mathrm{d}^{2}x\sqrt{-g}\quantity[(\sigma_{c}-\sigma_{b})R[g]+(\delta\varphi_{c}-\delta\varphi_{b})(R[g]-2)]+I_{\rm quadratic}\leavevmode\nobreak\ ,$ (30)
where
[…]
We can then eliminate $I_{\text{quadratic}}$ by requiring $\delta\varphi_{c}=-\delta\varphi_{b}$. Therefore, the scalar fluctuations in the UV brane locations are described by dS JT gravity, which indicates the configuration is stable after the orbifold projection. Notice that the theory will get increasingly suppressed as the UV brane approaches the AdS boundary.
@@ 3 Coarse-graining of information in the multiverse
In this section, we study the coarse-grained information that can be collected by a UV brane observer in the multiverse model in 2.2 using holographic entanglement entropy and compare it with the island formula in false vacuum decay models in quantum cosmology Aguilar-Gutierrez:2021bns ; Levine:2022wos and the cosmological central dogma Shaghoulian:2021cef . We find agreement with the previous literature, and new insights from the perspective of entanglement between the UV and IR energy modes in our model.
@@ 3.1 Information recovery protocol
Our protocol, based on Aguilar-Gutierrez:2023tic , considers an observer collecting Hawking radiation in a single double-sided brane $Q_{c}$, which is close enough to the asymptotic boundary (i.e. $\ell_{d+1}\rightarrow 0$ and $G_{d}\rightarrow 0$), such that the higher curvature corrections can be suppressed. The latter allows us to use the HRT formula as well as renders the bath non-gravitating. Gravity decouples and the effective theory of the UV brane is thus described by a dS QFT with a mass gap Maldacena:2012xp ; Fischler:2013fba . We work in the framework of IR/UV entanglement, where, as explained in Balasubramanian:2011wt ; Aguilar-Gutierrez:2023tic , the total Hilbert space factorizes as [^11] The argument employs a Fock space decomposition, $\mathcal{H}=\otimes_{\vec{p}}\mathcal{H}_{\vec{p}}$ with $\mathcal{H}_{\vec{p}}$ the Hilbert space for momentum modes $\vec{p}$.
@@
where $\mathcal{H}_{\text{UV}}$ is the Hilbert space of the dS QFT (with a gap) and $\mathcal{H}_{\text{IR}}$ represents the Hilbert space of the IR degrees of freedom, geometrized by $Q_{b}$. Also, note that, although there is no gravity localization in the IR brane, in contrast to the UV brane, this does not alter the evaluation of the HRT formula from the UV boundary.
@@
We now proceed with the evaluation in (33) within a single dS wedge holographic setting, meaning a single AdS3 space capped off by the UV and IR branes. We evaluate the HRT surfaces anchored at the UV brane, whose cutoff surface $\sigma_{c}\rightarrow\infty$ corresponds to $\rho_{c}\rightarrow\infty$.
Our interest is to dress the dS braneworlds with dS JT gravity to model false vacuum eternal inflation as in Aguilar-Gutierrez:2021bns (a perioidic dS2 spacetime with multiple inflationary and black hole patches). We have the option to either add intrinsic gravity on the UV and/or IR branes. In the first case, given that the UV brane represents a weakly gravitating region with Dirichlet boundary conditions, this choice would only amount to a shift in the entropy (43) both before and after the Page transition. Since we are only interested in the entropy difference between the two phases, which is unaffected by adding JT on the UV brane, we will only include JT on the IR brane, which has a nontrivial modification in the variational problem.
@@ 3.2 Before the Page transition
We proceed to determine the conserved charges from (34, 35) related to the angular momentum and energy in the ambient AdS3 space
@@
In the following, we show that an observer in a given (double-sided) UV braneworld will detect an increase in the von Neumann entropy until reaching a Page transition Aguilar-Gutierrez:2021bns .
@@ 3.3 Transition without JT couplings
$S^{\rm bef\,PT}=\frac{1}{4G_{3}}\quantity(\sigma_{c}-\log\absolutevalue{\sec\Delta\alpha})$ (44)
@@
3.4 DS JT gravity on the IR brane
We will now perform the extremization in (35) again, but now the conserved charge $E_{\alpha}$ does no longer vanish at the brane location. We start with the global AdS coordinates (3), with two conserved charges in $d=2$. One can then use the variation of the total action and the solutions to the EOM to evaluate the charge with Neumann boundary conditions. We find,
@@ 3.4 DS JT gravity on the IR brane
Imposing Dirichlet boundary conditions at the UV brane, the vanishing of the variation above gives us
@@
We may now evaluate the functional $S^{\rm aft\,PT}$ in (47) subject to (51, 52). However, we need to express $t_{b}$ in terms of the physical parameters measured by the UV observer, namely the angular distance for collecting the radiation $\alpha(\sigma_{c})$, and the physical time $t_{c}$. We will then have to solve the equations of motion of the extremal area surfaces and impose boundary conditions at $\sigma_{c}$, to determine those that minimize $S^{\rm after\leavevmode\nobreak\ PT}$.
@@
where $\Delta\tau=\tau(\rho)-\tau_{0}$. At the location of the UV brane, we get
@@
Let us now solve (56, 58) to determine two branches of solutions for $E_{\alpha}$ and $E_{\tau}$:
@@
Moreover, we have a relation between the charges above with the brane coordinates $t_{b}$, $\alpha_{b}$ in (51, 52), which is enough to determine the solution to the extremization problem.
@@
With all tools at hand, we can evaluate the island transition using (47, 48) as
@@
where $\epsilon_{c}$ is a UV regulator, and we have set $\varphi_{1}=0$. Since there are 2 roots in (59) and (60), we must explore the one that produces the minimum entropy to identify the island contribution. The comparison between the roots is displayed in Fig. 3, which shows that the dominating saddle is the one determined through $E_{\tau}^{(-)}$, $E_{\alpha}^{(-)}$ in (59, 60). Moreover, one can notice that the entropy for both curves starts decreasing until it reaches an asymptotic late time value, which we deduce below. The decrease in von Neumann entropy can then be used for information recovery using protocols such as Aalsma:2022swk .
@@
We now deduce the late-time expression for the entropy with the above relations. Using the ansatz (61, 62), one finds
@@
and therefore, the von Neumann entropy collected by a UV brane observer at $t_{c}\gg 1$ for generic $\Delta\alpha_{b}$ is just the constant given in (66).
@@
where $\epsilon_{c^{\prime}}$ is a UV regulator when $\sigma_{c}\rightarrow\infty$, and in the last expressing we have discarded $t_{c}$ independent additive terms. The plot of the Page curve, displaying the dominating island and the Hawking phase is shown in Fig. 4.141414The resulting island phase differs from other approaches in the literature to model multiverses from Karch-Randall branes Yadav:2023qfg , where instead, it was suggested that the total entropy should be the sum of the individual contribution from each brane and its bath. The difference with respect to our model is that we glue UV and IR branes together through Israel matching conditions, while the latter approach patches them all together in a codimension-two conformal defect.

### v3

@@ 2.3 Multiverse models from the (JT) dS wedge holography
We now describe a multiverse model by gluing n-copies of dS wedge universes. The different UV and IR regions are pairwise glued together in the configuration illustrated in Fig. 2.
Figure 2: Penrose diagram of the AdS3 space (in gray) cutoff by UV and IR dS regions (in red). We prepare the state with the Hartle-Hawking construction, considering a thermal density matrix for each brane $\rho_{\rm TFD}$ purified by a maximally entangled partner (UV or IR region respectively) with an empty bulk in between (white) and glue them together (red arrows), as shown in the figure. Although the UV and IR regions within a single universe are initially unentangled, as the branes accelerate and exchange Hawking modes.
@@
    $n$ double-sided IR and $n$ UV dSd regions with a matter theory entangled with each other.
@@
    $n$ double-sided AdSd+1 bulk space cutoff in the UV and IR region that glued along pairs of UV/UV or IR/IR regions periodically.
In the simplest case, $d=2$, gravity would be topological. To generate a dynamical evolution with respect to a braneworld observer, we add an intrinsic gravitational theory, namely dS JT gravity, on the IR brane. [^12] Alternatively, we could have added dS JT gravity on both the UV and IR regions. One can immediately see that adding JT on the UV region would modify the phases for the entanglement in (28, 29) by the same additive (time-dependent) constant. Since only the relative entropy difference enters in the physical analysis, we do not consider this possibility, although it can be straightforwardly included. We now analyze the conditions for this configuration to be realized.
@@
To set the stage, consider dS JT gravity as intrinsic theory on the IR branes, so that the system is described by
@@
The brane equation of motion translates to the Israel junction conditions Israel:1966rt
$\Delta K_{ij}-h_{ij}\Delta K=8\pi G_{3}T_{ij}$ (16)
where $T_{ij}$ is the effective stress tensor on the brane; while $\Delta K_{ij}=K_{ij}^{(+)}-K_{ij}^{(-)}=2K_{ij}^{(+)}$ is the extrinsic curvature difference between opposite sides, denoted by $\pm$, in the double-sided $Q_{b}$ brane.
On the other hand, the dilaton equation of motion fixes the background geometry on the ETW brane to be
@@
As last remark, note that unlike the original formulation of dS wedge holography in Aguilar-Gutierrez:2023tic , since we consider a single ETW brane in the bulk, there is no longer issue about perturbative stability due to fluctuations of the ETW brane location.
@@ 3 Coarse-graining of information in the multiverse
In this section, we study the coarse-grained information that can be collected by a UV observer in the multiverse model in 2.3 using holographic entanglement entropy and compare it with the island formula in false vacuum decay models in quantum cosmology Aguilar-Gutierrez:2021bns ; Levine:2022wos and the cosmological central dogma Shaghoulian:2021cef .
As mentioned in Sec. 1, we stress that having the UV cutoff very close to the asymptotic boundary allows for suppressing non-local effects on the dual QFT (where the scale of non-locality depends on the deformation parameter, as $\sqrt{\lambda}$ Guica:2019nzm ). The Ryu-Takayanagi (RT) formula Ryu:2006bv ; Ryu:2006ef is generically not applicable due to the non-locality of the field theory, but it remains well defined at the perturbative level in the deformation parameter $\lambda$ Banerjee:2024wtl , or at the non-perturbative one, considering that the deformation parameter is the smallest length scale in the system Apolo:2023ckr . For this reason, we place the UV cutoff very near the conformal boundary, meaning that we keep the deformation parameter $\lambda\ll 1$ (i.e. the cutoff location $\sigma_{c}\gg 1$ in (13)), where the theory accepts a quasilocal description, and the holographic entanglement entropy is well-defined Apolo:2023ckr ; Banerjee:2024wtl ; Demise:2021cfx ; Chen:2018eqk ; Donnelly:2018bef .
@@ 3.1 Information recovery protocol
Our protocol, based on Aguilar-Gutierrez:2023tic , considers an observer collecting Hawking radiation in a single double-sided cutoff region $Q_{c}$ with Dirichlet boundary conditions. Gravity decouples and the effective theory of the UV region is thus described by a dS QFT with a mass gap Maldacena:2012xp ; Fischler:2013fba . We work in the framework of IR/UV entanglement, where, as explained in Balasubramanian:2011wt ; Aguilar-Gutierrez:2023tic , the total Hilbert space factorizes as [^13] The argument employs a Fock space decomposition, $\mathcal{H}=\otimes_{\vec{p}}\mathcal{H}_{\vec{p}}$ with $\mathcal{H}_{\vec{p}}$ the Hilbert space for momentum modes $\vec{p}$, so that the notion of entanglement between UV and IR degrees of freedom in the dual theory are described at different momentum scales Balasubramanian:2011wt .
@@
where $\mathcal{H}_{\text{UV}}$ is the Hilbert space of the dS QFT (with a gap) and $\mathcal{H}_{\text{IR}}$ represents the Hilbert space of the IR degrees of freedom, geometrized by $Q_{b}$. Also, note that, although there is no gravity localization in the IR brane, this does not alter the evaluation of the HRT formula from the UV boundary.
@@
We now proceed with the evaluation in (27) within a single $T^{2}$ dS wedge holographic universe, meaning a single AdS3 space capped off by the UV and IR regions. We evaluate the HRT surfaces anchored at the finite UV region, whose cutoff surface $\sigma_{c}\rightarrow\infty$ corresponds to $\rho_{c}\rightarrow\infty$.
Our interest is to dress the dS braneworlds with dS JT gravity to model false vacuum eternal inflation as in Aguilar-Gutierrez:2021bns (a periodic dS2 spacetime with multiple inflationary and black hole patches). We have the option to either add intrinsic gravity on the UV and/or IR regions. In the first case, given that the UV region represents a region with Dirichlet boundary conditions, this choice would only amount to a shift in the entropy (37) both before and after the Page transition. Since we are only interested in the entropy difference between the two phases, which is unaffected by adding JT on the UV region, we will only include JT on the IR brane, which has a nontrivial modification in the variational problem.
@@ 3.2 Before the Page transition
We proceed to determine the conserved charges from (28, 29) related to the angular momentum and energy in the ambient AdS3 space
@@
In the following, we show that an observer in a given (double-sided) finite cutoff UV region will detect an increase in the von Neumann entropy until reaching a Page transition.
@@ 3.3 Transition without JT couplings
$S^{\rm bef\,PT}=\frac{1}{4G_{3}}\quantity(\sigma_{c}-\log\absolutevalue{\sec\Delta\alpha})\leavevmode\nobreak\ .$ (38)
@@
3.4 De Sitter JT gravity on the IR brane
We will now perform the extremization in (29) again, but now the conserved charge $E_{\alpha}$ no longer vanishes at the brane location. We start with the global AdS coordinates (3), with two conserved charges in $d=2$. One can then use the variation of the total action and the solutions to the EOM to evaluate the charge with Neumann boundary conditions. We find,
@@ 3.4 De Sitter JT gravity on the IR brane
Imposing Dirichlet boundary conditions at the finite UV cutoff region, the vanishing of the variation above gives us
@@
We may now evaluate the functional $S^{\rm aft\,PT}$ in (41) subject to (45, 46). However, we need to express $t_{b}$ in terms of the physical parameters measured by the UV observer, namely the angular distance for collecting the radiation $\alpha(\sigma_{c})$, and the physical time $t_{c}$. We will then have to solve the equations of motion of the extremal area surfaces and impose boundary conditions at $\sigma_{c}$, to determine those that minimize $S^{\rm after\leavevmode\nobreak\ PT}$.
@@
where $\Delta\tau=\tau(\rho)-\tau_{0}$. At the location of the UV region, we get
@@
Let us now solve (50, 52) to determine two branches of solutions for $E_{\alpha}$ and $E_{\tau}$:
@@
Moreover, we have a relation between the charges above with the brane coordinates $t_{b}$, $\alpha_{b}$ in (45, 46), which is enough to determine the solution to the extremization problem.
@@
With all tools at hand, we can evaluate the island transition using (41, 42) as
@@
where $\epsilon_{c}$ is a UV regulator, and we have set $\varphi_{1}=0$. Since there are two roots in (53) and (54), we must explore the one that produces the minimum entropy to identify the island contribution. The comparison between the roots is displayed in Fig. 3, which shows that the dominating saddle is the one determined through $E_{\tau}^{(-)}$, $E_{\alpha}^{(-)}$ in (53, 54). Moreover, one can notice that the entropy for both curves starts decreasing until it reaches an asymptotic late time value, which we deduce below. The decrease in von Neumann entropy can then be used for information recovery using protocols such as Aalsma:2022swk .
@@
We now deduce the late-time expression for the entropy with the above relations. Using the ansatz (55, 56), one finds
@@
and therefore, the von Neumann entropy collected by a UV observer at $t_{c}\gg 1$ for generic $\Delta\alpha_{b}$ is just the constant given in (60).
@@
where $\epsilon_{c^{\prime}}$ is a UV regulator when $\sigma_{c}\rightarrow\infty$, and in the last expression we have discarded $t_{c}$ independent additive terms. The plot of the Page curve, displaying the dominating island and the Hawking phase is shown in Fig. 4.161616The resulting island phase differs from other approaches in the literature to model multiverses from Karch-Randall branes Yadav:2023qfg , where instead, it was suggested that the total entropy should be the sum of the individual contribution from each brane and its bath. The difference with respect to our model is that the UV and IR regions intercept only at two boundary times (see Fig. 1), while the latter approach patches them all together in a codimension-two conformal defect.